OpenStudy (anonymous):

Can someone help me with this problem?. Ri(t)+Ldi(t)dt+1/C∫i(t)dt=E(t)

OpenStudy (anonymous):

(1) I have this equation (RLC circuit): \[Ri(t)+L \frac{ di(t) }{ dt }+\frac{ 1 }{ C }\int\limits_{?}^{?}i(t)dt=E(t)\] We use this: \[x_{1}(t)=\int\limits_{?}^{?}i(t)dt\] \[x_{2}=\frac{ dx_1{1}(t) }{ dt }=i(t)\] (2) Then we substitute: \[R \frac{ dx_{1}(t) }{ dt }+L \frac{ d^{2}x_{1}(t) }{ dt^{2} }+\frac{ 1 }{ C }x_{1}(t)=E(t)\] And : \[x_{2}=\frac{ dx_1{1}(t) }{ dt }=i(t)\], so \[\frac{ di(t) }{ dt }=\frac{ d^{2}x_{1}(t) }{ dt^{2} }=\frac{ dx_{2}(t) }{ dt }\] .My question is how can I decompose the (2) in differential equations of first order?. Because I didn't see that in the last semester. Can you explain me with a different example, because I want to do this by my own. Thanks One of them is: \[\frac{ dx_{1} }{ dt }=x_{2}(t)\]. But I don't know why

OpenStudy (anonymous):

I dont really understand what is that you are asking for. In this case \(x_1(t)\) is the charge Q stored in time interval \(t_1-t_0\). So dQ/dt=i(t)

OpenStudy (anonymous):

This is a sentences: A differential equation of n-th order can be decomposed into n-th differential equations of first order. But I don't know how

OpenStudy (anonymous):

ok. I will try to explain...

OpenStudy (anonymous):

imagine you have 2º order ODE: y''=f(x,y,y') then you can introduce 2 new variables: u=y and v=y' Using this, we see that u'=v, right till here?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so, continuing v'=y''=f(x,y,y') your original equation. So you got two equations: 1º: u'=v 2º v'=f(x,u,v) which represent same original relation

OpenStudy (anonymous):

Ahh, ok.

OpenStudy (anonymous):

I think I got it thanks.

OpenStudy (anonymous):

How it is called this method in English?